### v1

@@ 3.1 Predictions of Hurricane Florence
In contrast, FourCastNetV2 exhibits superior performance throughout the entire lifecycle of the hurricane (see Fig. 3 (b)). The prediction remains accurate even after landfall, suggesting that FourCastNetV2 is better equipped to handle the complexities of land interaction. This improvement is explained by advancements in model architecture, the advent of the Spherical Fourier Neural Operator (Bonev 2023) and the more robust training FourCastNetv2 underwent compared to FourCastNet.
Following the tutorial guidance we have provided in the GitHub repository (Khadir and Stevenson 2025), a student can make the same predictions and reproduce Fig. 3. This demonstrates what undergraduate students and low-resource research groups can achieve when given guided access to powerful AI tools, such as FourCastNet and FourCastNetv2. Here, the computer code for data visualization and the model run were entirely made by undergraduate student researchers with references to the existing climate data visualization code in modern textbooks and their online materials, e.g., Shen and North (2023) and Shen and Somerville (2019). The visualization code and guidance to access the correct data are included in our user’s manual in GitHub (Khadir and Stevenson 2025).
3.2 Predictions of Zonal Winds and Surface Pressure
@@
Figure 5 presents two predictions of surface pressure from FourCastNetv2 at 24 hours and 150 hours from the initial time Z00:00, 25 December 2024.
@@ 3.3 Setup and Results for Training FourCastNet
Thus far, we have primarily explored the power of inference, demonstrating how FourCastNet and FourCastNetv2 can be leveraged by undergraduates and resource-limited research groups for meaningful contributions to atmospheric research. Another important aspect of utilizing these AI models is training. Next, we present the setup and results of training FourCastNet.
@@
Model training requires substantial computational resources, particularly for large-scale AI-driven weather forecasting. The original FourCastNet model was trained on a cluster of 64 A100 GPUs over 16 hours (Pathak et al. 2022), a setup far beyond the reach of many research institutions. However, advancements in computing accessibility have enabled universities and organizations to allocate powerful hardware to their students. San Diego State University’s NRP Hyper Cluster is one such example, providing access to high-performance GPUs for training deep-learning models.
@@
Using four A100 GPUs from the NRP Hyper Cluster, we fine-tuned FourCastNet’s backbone for an additional 18 epochs. This training process spanned 110 hours with a batch size of 16, significantly smaller than NVIDIA’s original batch size of 64. To adapt the model to our system, we modified several scripts and configurations from the original FourCastNet release. A complete guide to reproducing our training setup, including necessary configurations and environment details, is available in our user’s manual on GitHub (Khadir and Stevenson 2025).
@@
In "FourCastNeXt: Optimizing FourCastNet Training for Limited Compute" (Guo 2024), the authors achieved improved results only when they began training on a different, more diverse dataset, which allowed FourCastNet to continue learning and refining its predictions. Our experience confirmed these findings when, throughout all 18 epochs, training and validation loss oscillated around the same value. This plateau was not due to insufficient training effort but was a result of the model having already converged on the provided data and weights. Essentially, without introducing new data for further learning, the model was unable to make additional improvements. This highlights the importance of having access to fresh and varied data to drive further improvements in model performance.
@@
The ability to connect our A100 GPUs to FourCastNet’s complex system for AI model training is another significant step. A major barrier for undergraduate students and resource-limited research groups is the technical knowledge required to set up and effectively utilize advanced GPU systems. Integrating one system with another to enable the use of AI models for training is often the most challenging task in the entire AI weather forecasting process. For example, the FourCastNet scripts were designed for a Simple Linux Utility for Resource Management (SLURM) based system, while the system we had access to was Kubernetes-based. This mismatch required a deep understanding of both systems and the ability to translate the necessary configurations and scripts from one environment to another. The technical expertise to adapt these resources and integrate them seamlessly is a key capability for researchers who wish to leverage AI for atmospheric research. The close collaboration of San Diego State University’s Research and Cyberinfrastructure team allowed us to interact with the necessary resources through a Kubernetes environment specially created with the appropriate packages installed to use FourCastNet (see Figs. 1 and 2 about the setup of the collaboration and hardware). By successfully configuring and adapting FourCastNet to work with our available hardware, we have demonstrated that even researchers with limited expertise in GPU systems can still leverage high-performance resources for deep-learning tasks. Ultimately, our goal is to lower the entry barriers and empower a broader community of researchers to explore and utilize AI-driven models, which is crucial for advancing AI in fields like weather forecasting.
@@ 5 User’s Manual to Make Weather Forecasts Using FourCastNetv2
Users of our GitHub repository https://github.com/ikhadir/FourCastNet-with-JupyterLabs should make adjustments to our sample code where necessary. Options for visualizations of the data and predictions are available to suit the prediction being created. We visualize the data using a Pacific-centered Robinson projection with grid lines marking longitude and latitude. The color map that is selected can be determined for the specific variable to be shown. We encourage users to treat the manual as a tool to explore the AI forecasting model in its entirety. We have also created tools to improve the visualizations of the raw data, global predictions, and hurricane tracking. These tools are straightforward to use and allow for a multitude of configurations and adjustments to fit the user’s preferences and demands.
@@ 6 Discussion and Conclusions
The existing fast and convenient AI forecasting work will likely lead to long-term progress in democratizing AI models and, hence, give students hands-on experience in AI weather predictions. Among numerous opportunities, the following three areas may present immediate and practical applications for future research exploration: (a) AI weather predictions for small island nations, (b) training students in classrooms using FourCastNet, and (c) generating predictions with real-time satellite data.
@@
For area (a), we can envision applications of the AI weather forecasting model around the world to solve a variety of problems. One such application is the predictive power of the model for remote islands. The experience of North Carolina State University meteorology students provided crucial wind forecasts for a team of ecologists on Bouvet Island (Riordan 2003). Communication and data traffic to the island were limited due to the bandwidth constraint. Weather predictions were crucial for the completion of the research as well as the safety of the crew, one of whom was already injured by strong winds. Riordan describes the wind predictions his team provided as having marginal reliability at a 48-hour lead time. Running FourCastNetv2 locally will avoid the problem of slow data transmission and allow island residents to create their own forecasts and act accordingly. These forecasts will be valuable to researchers and local populations in preparing for agricultural needs, preventing flash flood damage, and taking appropriate measures for public safety during hazardous weather events.
As an example of this, we provide the U10 predictions for the Pacific Island nation of Vanuatu, which is particularly vulnerable to tropical cyclones (see Fig. 8).
@@
For area (b), in this paper, we wish to share our student experience with the FourCastNet AI weather models. The audience of this paper will find that they benefit better from our experience if they are more familiar with some of the technical details of FourCastNetv2, such as the variables which the model uses and the output file which the predictions produce. The first two authors of this paper are undergraduate students and encountered certain challenges that could be avoided by other university students who implement the model. For a successful implementation, students will need some familiarity with Bash to interact with the local development environment and coordinate hardware configurations with the model itself. As indicated in Figure 1, strong support from a dedicated IT team in a university is crucial for setting up the necessary software and hardware configurations. The SDSU Climate Informatics Lab (SCIL) has been instrumental in providing consultation for this project. It is essential that students have the backing of a supportive principal investigator, such as the SCIL director, Samuel Shen, in our case. It is also important to have a committed team of lab members who collaboratively ensure that the AI weather forecasting research meets high-quality and professional standards.
@@
For area (c), some university groups may aggregate satellite data directly without a data assimilation step and use the aggregated data on a 4D space-time grid to allow the initiation of the AI weather forecasting model to perform near real-time forecasts. Thus far, our AI forecastings are limited to using the assimilated ERA5 output as initialization. Although ERA5 data can be pulled using API, there is a five-day delay for the present data to be uploaded. Using the space-time gridded satellite data directly will allow for near real-time forecasts to be produced.
Our analysis of FourCastNet and FourCastNetv2 demonstrates the potential of advanced AI models for weather prediction, particularly in the context of limited computational resources faced by university research groups. NVIDIA’s FourCastNet models drastically reduce the time and cost of creating weather forecasts compared to traditional Numerical Weather Prediction (NWP) methods, but there are still a few barriers for small research groups to replicate AI weather forecasting results. Our work intends to resolve some practical challenges of using GPUs, such as NVIDIA A100, for large-scale machine learning applications in numerical weather forecasting.
7 Data and Code Availability
The tutorial, computer code, and instructions for this paper are freely available at https://github.com/ikhadir/FourCastNet-with-JupyterLabs. FourCastNet from NVIDIA is openly accessible at https://github.com/NVlabs/FourCastNet. FourCastNetv2 can be accessed using the ECMWF command line plug-in with https://github.com/ecmwf-lab/ai-models. The ERA5 data set can be accessed through an account from the Climate Data Store, the full set is about 6 TB.
@@ 8 Acknowledgments
We thank the collaboration and support from the SDSU Climate Informatics Laboratory, in particular, the indispensable guidance of Danielle Lafarga. The Research and Cyberinfrastructure group at SDSU provided constant technical support for hardware setup and software assimilation. This research used resources from the U.S. National Science Foundation (NSF)-funded Technology Infrastructure for Data Exploration (TIDE) project (Award No. 2346701) and from the the National Research Platform (NRP) at the University of California, San Diego (NSF award No. 1730158, 1540112, 1541349, 1826967, 2112167, 2100237, and 2120019). This work was also partially supported by a U.S. NSF grant (Award No. 2324008) and a U.S. NOAA grant (Award No. NA22SEC4810016). We thank researchers at ECMWF for maintaining the ERA5 dataset. Sincere thanks to Jacob Radford from the University of Colorado for his guidance in running the AI-models plug-in. We thank Mitch Goldberg’s group at City University of New York for the opportunity to demonstrate our work. Finally, we thank Amy McGovern, Director of the U.S. National Science Foundation AI2ES project (Award No. 2019758), for bridging the collaboration between the SDSU and NVIDIA researchers on this work.

### v3

@@ 2.1.2 Prediction Results
In contrast, FourCastNetV2 exhibits superior performance throughout the entire lifecycle of the hurricane (see Fig. 3 (b)). The prediction remains accurate even after landfall, suggesting that FourCastNetV2 is better equipped to handle the complexities of land interaction. This improvement is explained by advancements in model architecture, the advent of the Spherical Fourier Neural Operator (Bonev et al., 2023) and the more robust training FourCastNetv2 underwent compared to FourCastNet.
Following the tutorial guidance we have provided in the GitHub repository (Khadir and Stevenson, 2025), a student can make the same predictions and reproduce Fig. 3. This demonstrates what undergraduate students and low resource research groups can achieve when given guided access to powerful AI tools, such as FourCastNet and FourCastNetv2. Here, the computer code for data visualization and the model run were entirely made by an undergraduate student research with references to the existing climate data visualization code in modern textbooks and their online materials, e.g., Shen and North (2023) and Shen and Somerville (2019). The visualization code and guidance to access the correct data is included in our user’s manual in GitHub (Khadir and Stevenson, 2025).
2.2 Global Predictions
2.2.1 Zonal Wind Predictions
@@
2.2.2 Surface Pressure Predictions
Figure 5 presents two predictions of surface pressure from FourCastNetv2 at 24 hours and 150 hours from the initial time Z00:00 25 December 2024.
@@ 2.3 Setup and Results for Training FourCastNet
Thus far, we have primarily explored the power of inference, demonstrating how FourCastNet and FourCastNetv2 can be leveraged by undergraduates and resource-limited research groups for meaningful contributions to atmospheric research. Another important aspect of utilizing these AI models is training. In this paper, we use FourCastNet as an example to describe the setup and results of training.
@@
Model training requires substantial computational resources, particularly for large-scale AI-driven weather forecasting. The original FourCastNet model was trained on a cluster of 64 A100 GPUs over 16 hours (Pathak et al., 2022), a setup far beyond the reach of many research institutions. However, advancements in computing accessibility have enabled universities and organizations to allocate powerful hardware to their students. San Diego State University’s NRP Hyper Cluster is one such example, providing access to high-performance GPUs for training deep-learning models.
@@
Using four A100 GPUs from the NRP Hyper Cluster, we fine-tuned FourCastNet’s backbone for an additional 18 epochs, spanning 110 hours of training with a batch size of 16, significantly smaller than NVIDIA’s original batch size of 64. To adapt the model to our system, we modified several scripts and configurations from the original FourCastNet release. The model was trained on ERA-5 global data from 1979 to 2015, with the years 2016 and 2017 reserved for testing, and 2018 held out as a dedicated validation set to assess its generalization performance on unseen data. Figure 6 shows validation loss across the 18 training epochs when training FourCastNet. A complete guide to reproducing our training setup, including the necessary configurations and environment details, is available in our user manual on GitHub (Khadir and Stevenson, 2025).
@@
In "FourCastNeXt: Optimizing FourCastNet Training for Limited Compute" (Guo et al., 2024), the authors found that improved results were only achieved when training on a more diverse dataset, allowing FourCastNet to continue refining its predictions. Our experience confirmed this, as throughout all 18 epochs, both the training and validation loss oscillated around the same value. We theorize that this plateau was not due to insufficient training effort, but because the model had already converged on the provided data and weights. Figure 6 illustrates the validation loss for FourCastNet across these 18 epochs, where one epoch represents a pass through the entire dataset (ERA-5 1979 to 2015). The points show the mean squared error (MSE) between FourCastNet’s predictions and the ground truth of the validation dataset. The stagnation of the model’s validation loss indicates that it was unable to improve on the current dataset.
@@
The ability to connect our A100 GPUs to FourCastNet’s complex system for AI model training is another significant step. A major barrier for undergraduate students and resource-limited research groups is the technical knowledge required to set up and effectively utilize advanced GPU systems. Integrating one system with another to enable the use of AI models for training is often the most challenging task in the entire AI weather forecasting process. For example, the FourCastNet scripts were designed for a Simple Linux Utility for Resource Management (SLURM) based system, while the system we had access to was Kubernetes based. This mismatch required a deep understanding of both systems and the ability to translate the necessary configurations and scripts from one environment to another. The technical expertise to adapt these resources and integrate them seamlessly is a key capability for researchers who wish to leverage AI for atmospheric research. The close collaboration of San Diego State University’s Research and Cyberinfrastructure team allowed for us to interact with the necessary resources through a Kubernetes environment specially created with the appropriate packages installed to use FourCastNet (see Figs. 1 and 2 about the setup of the collaboration and hardware). By successfully configuring and adapting FourCastNet to work with our available hardware, we have demonstrated that even researchers with limited expertise in GPU systems can still leverage high-performance resources for deep-learning tasks. Ultimately our goal is to lower the entry barriers and empower a broader community of researchers to explore and utilize AI-driven models, which is crucial for advancing AI in fields like weather forecasting.
@@ 4 User’s Manual to Make Weather Forecasts Using FourCastNetv2
Users of our GitHub repository https://github.com/ikhadir/FourCastNet-with-JupyterLabs should make adjustments to our sample code where necessary. Options for visualizations of the data and predictions are available to suit the prediction being created. We visualize the data using a Pacific-centered Robinson projections with grid lines marking longitude and latitude. The color map that is selected can be determined for the specific variable to be shown. We encourage users to tread the manual as a tool to explore the AI forecasting model in its entirety. We have also created tools to improve the visualizations of the raw data, global predictions, and hurricane tracking. These tools are straightforward to use and allow for a multitude of configurations and adjustments to fit the user’s preferences and demands.
@@ 5 Discussion and Conclusions
The existing fast and convenient AI forecasting work will likely lead to a long term progress to democratize AI models and hence give students hands on experience in AI weather predictions. Among numerous opportunities, the following three areas may present immediate and practical applications for future research exploration: (a) AI weather predictions for small island nations, (b) training students in classrooms using FourCastNet, and (c) generating predictions with real-time satellite data.
@@
For area (a), we can envision applications of the AI weather forecasting model around the world to solve a variety of problems. One such application is the predictive power of the model for remote islands. The experience of North Carolina State University meteorology students provided crucial wind forecasts for a team of ecologists on Bouvet Island (Riordan, 2003). Communication and data traffic to the island were limited due to the bandwidth constraint. Weather predictions were crucial for the completion of the research as well as the safety of the crew, one of whom was already injured by strong winds. Riordan describes the wind predictions his team provided as having marginal reliability at 48-hour lead time. Running FourCastNetv2 locally will avoid the problem of slow data transmission and allow island residents to create their own forecasts and act accordingly. These forecasts will be valuable to researchers and local populations in preparing for agricultural needs, preventing flash flood damage, and taking appropriate measures for public safety during hazardous weather events. As an example of this, we provide the U10 predictions for the Pacific Island nation of Vanuatu, which is particularly vulnerable to tropical cyclones (see Fig. 9).
@@
For area (b), in this paper, we wish to share our student experience with the FourCastNet AI weather models. The audience of this paper will find that they benefit better from our experience if they are more familiar with some of the technical details of FourCastNetv2, such as the variables which the model uses and the output file which the predictions produce. The first two authors of this paper are undergraduate students and encountered certain challenges that could be avoided by other university students who implement the model. For a successful implementation, students will need some familiarity with Bash to interact with the local development environment and coordinate hardware configurations with the model itself. As indicated in Figure 1, strong support from a dedicated IT team in a university is crucial for setting up the necessary software and hardware configurations. The SDSU Climate Informatics Lab (SCIL) has been instrumental in providing consultation for this project. It is essential that students have the backing of a supportive principal investigator, such as the SCIL director Samuel Shen in our case. It is also important to have a committed team of lab members who collaboratively ensure that the AI weather forecasting research meets high-quality and professional standards.
@@
For area (c), some university groups may aggregate satellite data directly without a data assimilation step and use the aggregated data on a 4D space-time grid to allow the initiation of the AI weather forecasting model to perform near real-time forecasts. Thus far, our AI forecastings are limited to use the assimilated ERA5 output as initialization. Although ERA5 data can be pulled using API, there is a five day delay for the present data to be uploaded. Using the space-time gridded satellite data directly will allow for near real time forecasts to be produced.
Our analysis of FourCastNet and FourCastNetv2 demonstrates the potential of advanced AI models for weather prediction, particularly in the context of limited computational resources faced by university research groups. NVIDIA’s FourCastNet models drastically reduce the time and cost of creating weather forecasts compared to traditional Numerical Weather Prediction (NWP) methods, but there are still few barriers for small research groups to replicate AI weather forecasting results. Our work intends to resolve some the practical challenges of using GPUs, such as NVIDIA A100, for large-scale machine learning applications in numerical weather forecasting.
6 Code and Data Availability
The initial version of FourCastNet is available under the NVIDIA Source Code License at https://github.com/NVlabs/FourCastNet under DOI: https://doi.org/10.5281/zenodo.16389569. FourCastNetv2 is the second version, accessible through the ECMWF AI Models Plugin at https://github.com/ecmwf-lab/ai-models under DOI: https://doi.org/10.5281/zenodo.16389730, released under the Apache License 2.0. A tutorial for using the initial version of FourCastNet and FourCastNetv2 is freely available at https://github.com/ikhadir/FourCastNet-withJupyterLabs under DOI: https://doi.org/10.5281/zenodo.16118616. The computer code for the tutorial used in this paper is also available at this DOI. We used the following NVIDIA materials: (i) The trained model weights, (ii) the preprocessing scripts, (iii) the ERA5-derived input data. These materials are publicly available via Globus at https://app.globus.org/file-manager?origin_id=945b3c9e-0f8c-11ed-8daf-9f359c660fbd&origin_path=%2F˜%2Fdata%2F. We used these materials to produce the results in this paper. The ERA5 reanalysis dataset used for the FourCastNetv2 model training and validation is available from the Copernicus Climate Data Store (CDS) at https://cds.climate.copernicus.eu upon account registration. We downloaded approximately 6TB of input data from NVIDIA’s Globus folder for our study in this paper.
7 Author contribution
IK, SS, and AB developed the Jupyter notebook tutorial; HL and KK created the relevant cyberinfrastructure to connect to SDSU GPU resources; IK, SS, and SSPS wrote the manuscript draft; HL, KK, DH, and SP reviewed and edited the manuscript. All coauthors have contributed to the original ideas and research objectives equally.
8 Competing interests
The authors declare that they have no conflict of interest.
@@ 9 Acknowledgments
We thank the collaboration and support from the colleagues at SDSU Climate Informatics Laboratory, in particular, the indispensable guidance of Danielle Lafarga. The Research and Cyberinfrastructure group at SDSU provided constant technical support for hardware setup and software assimilation. We thank researchers at ECMWF for maintaining the ERA5 dataset. Sincere thanks to Jacob Radford from the University of Colorado for his guidance in running the ai-models plug-in. We thank Mitch Goldberg’s group at City University of New York for the opportunity to demonstrate our work. Finally, we thank Amy McGovern, Director the U.S. National Science Foundation AI2ES project (Award No. 2019758), for bridging the collaboration between the SDSU and NVIDIA researchers on this work.
10 Financial Support
This research used resources from the U.S. National Science Foundation (NSF)-funded Technology Infrastructure for Data Exploration (TIDE) project (Award No. 2346701) and from the the National Research Platform (NRP) at the University of California, San Diego (NSF award No. 1730158, 1540112, 1541349, 1826967, 2112167, 2100237, and 2120019). This work was also partially supported by a U.S. NSF grant (Award No. IIS 2324008) and U.S. NOAA grants (Award No. NA22SEC4810016 and NA24NESX405C0006-T1-01).